OpenStudy (anonymous):

Find the derivative. y=3e^(-3/t)

OpenStudy (anonymous):

Let \(u = -3/t\).

OpenStudy (anonymous):

\[ y = 3e^{u} \]

OpenStudy (anonymous):

\[ \frac{dy}{dx} = \frac{dy}{du}\frac{du}{dx} \]

OpenStudy (anonymous):

Can you find these two derivatives?

OpenStudy (anonymous):

there is no u substitutions when finding the derivative

OpenStudy (anonymous):

What is \(dy/du\)?

OpenStudy (anonymous):

integrals

OpenStudy (anonymous):

I am not doing an integral.

OpenStudy (anonymous):

\[ y = 3e^u \]What is the derivative?

OpenStudy (anonymous):

Im not sure with that u

OpenStudy (anonymous):

Well, I want \(dy/du\) so you are differentiating with respect to \(u\). It is an easy one.

OpenStudy (anonymous):

Ok, can you walk me through this?

OpenStudy (anonymous):

Okay. Well \[ y=3e^{u}\implies \frac{dy}{du}=3e^u \]Does this make sense?

OpenStudy (anonymous):

so u=-3t^-1?

OpenStudy (anonymous):

Yes, just go with me for a moment. \[ u = -\frac{3}{t} \]

OpenStudy (anonymous):

We need to find \(du/dt\) now.

OpenStudy (anonymous):

du=3t^-2

OpenStudy (anonymous):

Ok, so now we say: \[ \frac{dy}{dt} = \frac{dy}{du}\frac{du}{dt} = 3e^u\frac{3}{t^2} \]Now, we can change \(u\) back into terms of \(t\): \[ \frac{dy}{dt}=3e^{-3/t}\frac{3}{t^2} \]

OpenStudy (anonymous):

Does this make sense?

OpenStudy (kainui):

@Bosse15 Trust wio, he's just doing the Chain Rule, it just looks a little different than what you're probably used to.

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

\[\int\limits_{?}^{?}xe ^{^{^{-3x ^{2}}}}dx\]

OpenStudy (anonymous):

Help with what?

OpenStudy (anonymous):

This time let \(u = -3x^2\)

OpenStudy (anonymous):

\[ du = d(-3x^2) =(-3x^2)'dx = -6xdx \]

OpenStudy (anonymous):

ok, then du=-6xdx

OpenStudy (anonymous):

-1/6 e^-3x^2+C?

OpenStudy (anonymous):

\[ \int e^{u}xdx \]We can do this: \[ =\frac{-6}{-6}\int e^{u}xdx = -\frac{1}{6}\int e^{u}(-6x)dx \]

OpenStudy (anonymous):

Then it becomes: \[ -\frac{1}{6}\int e^u\;du \]

OpenStudy (anonymous):

Oh, your answer is correct.

OpenStudy (anonymous):

Cool! lol thanks, could you stay on here just to make sure i can finish the rest of these up?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ e ^{\frac{ 1 }{ x }} }{ x ^{2} }\]

OpenStudy (anonymous):

dx\

OpenStudy (anonymous):

try using wolfram alpha

OpenStudy (anonymous):

never heard of that

OpenStudy (kainui):

www.wolframalpha.com It's really helpful for checking the answers to calculus problems.

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ e ^{2x} }{ e ^{2x}+1 }dx\]

OpenStudy (anonymous):

@wio

OpenStudy (kainui):

Wait I'm overcomplicating this. This is really quite a simple u-subsitution problem. e^2x+1=u